OpenStudy (anonymous):

How many solutions does the equation −3y + 3y + 4 = 4 have? Zero Infinitely many One Four

OpenStudy (anonymous):

Are you sure thats the right equatioon

OpenStudy (anonymous):

How many solutions does the equation −3y + 3y + 4 = 4 have? Zero Infinitely many One Four

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

okay..well i think the answer is 1, but im not possitive

geerky42 (geerky42):

Here, what is \(-3y + 3y\)? Try combine like terms

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Variable g is 8 more than variable w. Variable g is also 4 less than w. Which pair of equations best models the relationship between g and w? g = w − 8 g = w + 4 w = 8g w = g − 4 g = w + 8 g = w − 4 w = 8g w = g + 4

OpenStudy (lynfran):

g=w+4